# Common origin of black holes in high mass X–ray binaries and in gravitational–wave sources

### v3

@@
Common origin of black holes in high mass X–ray binaries and in gravitational–wave sources
Black hole (BH) high mass X-ray binary (HMXB) systems are likely the progenitors of the BH-BH mergers detected in gravitational waves by LIGO/Virgo/KAGRA (LVK). Yet merging BHs reach higher masses () than BHs in HMXBs () and have typically lower spins ( with a larger values tail) than often claimed for BH in HMXBs (). This could suggest that these two classes of systems belong to different populations, but we show here that this does not have to be the case. The difference in masses is easily explained as the known HMXB-BHs are in galaxies with relatively high metallicity, so their progenitor stars are subject to strong mass loss from winds, leading to relatively low mass BH at core collapse. Conversely, LVK can also detect BHs from low-metallicity galaxies that naturally produce more massive stellar-origin BHs. However, a difference in spin is more difficult to explain. Models with efficient angular momentum transport in stellar interiors produces slowly spinning progenitors of both LVK and HMXB BHs. Known HMXBs have orbital periods too long for efficient tidal spin-up, and are also unlikely to have undergone significant accretion spin-up. Instead we show that the derived value of the BH spin depends strongly on how the HMXB accretion disc emission is modelled. We argue that since Cyg X-1 is never observed to be in a soft spectral state, the appropriate spectral models must take into account the Comptonization of the disc photosphere. We show that such models are consistent with low spin values: . This has been confirmed recently by other teams for both Cyg X-1 and LMC X-1, and we show this is also the case for M33 X-7. We conclude that all known HMXB BHs can have low spin, in accordance with the results of stellar evolution models. Hence, current observations are consistent with LVK BHs and HMXB BHs belonging to the same population.
@@ 1 Introduction
LIGO/Virgo/KAGRA (LVK) interferometers have detected gravitational waves from double black hole (BH-BH) mergers (The LIGO Scientific Collaboration et al. 2023). These merging black holes (BHs) have masses in the range , with many primary (the more massive) BHs having masses of and (Callister & Farr 2024). The most massive LVK event (GW190521) showed two merging BHs with estimated mass of and . The majority of these mergers have low positive effective spin parameters, with the distribution that peaks at
@@
where denotes BH masses, are BH spin magnitudes (: BH angular momentum (AM), : the speed of light and the gravitational constant), and are angles between the system’s angular momentum and BH spins. Out of these mergers show high positive effective spins (mean), while none show high negative spin . However, there are events with moderate negative spin estimates: . Low values of the effective spin parameter may result from small individual BH spin magnitudes (Belczynski et al. 2020b). This finds support in LVK data that shows that BH individual spin magnitudes in BH-BH mergers peak at , although distribution has a long tail; that extends to large values (The LIGO Scientific Collaboration et al. 2023; see Callister & Farr 2024 for slightly different results).
@@
Electromagnetic observations have revealed a population of BHs hosted in binary star systems (e.g., see https://universeathome.pl/universe/blackholes.php and references therein). Out of several known binary configurations hosting BHs, only high-mass X-ray binaries (HMXBs) can potentially lead to BH-BH mergers, as BH companion stars are massive enough to form a BH.There are only three HMXBs known with such massive () companions. These have moderate-mass BHs: LMC X-1 with a star and a BH (Orosz et al. 2009), Cyg X-1 with a star and a BH (Miller-Jones et al. 2021) and M33 X-7 with a star and a BH (Ramachandran et al. 2022 revised down from the previous estimates of a donor and BH in Orosz et al. 2007). (We note that the binary LMC X-3 is not a HMBXB since its stellar companion is less massive () than the BH () Orosz et al. (2014) and won’t form a BH anyway.) Spins of BHs in HMXBs are obtained with two methods namely disc continuum fitting, and reflection spectroscopy. Reflection spectroscopy is complex for soft state spectra (e.g. Tomsick et al. 2018), so here we focus on spin values from disc continuum fitting. These are typically very high111The spins attributed through various methods to LMXB BH vary from very low to very high values, sometimes for the same system (e.g. Reynolds 2021). However, on the one hand these binary systems are not BHBH binary progenitors, on the other, their evolutionary history is still a puzzle (Wiktorowicz et al. 2014).:
@@
Since masses and spins of LVK BHs and HMXB BHs appear to be different, it might seem that these two groups of systems belong to populations with different formation scenarios. Here, we show that all these BHs may instead form a single population. The difference in BH mass arises naturally from the diverse formation environments: all known HMXB-BH are in local group galaxies with relatively high metallicity, so they have strong stellar winds before the core collapse, resulting in lower black hole masses compared to those forming in low metallicity galaxies which are within the sensitivity range of LVK. We demonstrate the difference in expected black hole masses explicitly using a population synthesis model. This model also reproduces the spin values in the LVK BHs, for models where high spins are produced only by tidal spin-ups in extremely close binaries. The known HMXB-BH systems are in binary systems which are not close enough for tidal spin-up. Hence the population synthesis models predict they should not have high spins, which conflicts with the reported high values. However, the disc continuum method relies on the emission thermalizing. If instead the disc is covered by a warm, optically thick skin, so that the emission is Comptonised, the spin values are all consistent with (Belczynski et al. 2021; Zdziarski et al. 2024a, b; this paper). With this interpretation, the LVK BH-BH mergers and the HMXBs can be members of the same binary star population.
@@ 2 BH Masses
To demonstrate the unicity of the BH population, we use the population synthesis code StarTrack (Belczynski et al. 2002, 2008) and employ the delayed core-collapse supernova (SN) engine for neutron star/BH mass calculation (Fryer et al. 2012) that does not produce a lower mass gap between NSs and BHs. We use a model in which the pair-instability (upper) mass gap is calculated with non-standard fusion reaction rates and heavy mixing for massive stars that allows for BHs to form up to (Farmer et al. 2020; Costa et al. 2021). Such model was already developed by Belczynski (2020) in the context of the detected at the low metallicity GW190521 LVK merger detected at distance of Gpc Abbott et al. (2020)).
In Figure 1 we show the dependence of the Helium core mass on metallicity for single stars obtained with the stellar wind treatment presented in Belczynski et al. (2010a). The helium core mass is a good approximation of the BH mass since binary interactions (Roche Lobe Overflow or common envelope evolution) tend to remove the H-rich envelope from stars in close binaries (progenitors of BH-BH mergers if these form through isolated binary evolution). The lower the metallicity, the lower the opacity in stellar atmospheres, and the lower the wind mass loss, which produces progressively higher mass BHs with decreasing metallicity.
Cyg X-1 is located in the Milky Way disc that hosts stars with about solar metallicity (, Asplund et al. (2009)). For such metallicity, we obtain a maximum BH mass of , which is consistent with the highest mass stellar-origin BH known in the Milky Way for the same Cyg X-1 ( Miller-Jones et al. (2021)). LMC X-1 is located in the Large Magellanic Cloud () and M33 X-7 in M33 galaxy (). In this metallicity range, BHs are not expected to be more massive than .
The LVK BH-BH mergers have already been detected (O3) to significant redshifts of . The detected BH-BH mergers may have formed even at much larger redshifts due to non-zero delay time between star formation and BH-BH mergers (Dominik et al. 2012; Fishbach & Kalogera 2021). Therefore, galaxies and stars that can produce BH-BH mergers cover a wide range of metallicity. It is even claimed that Population III (metal-free) stars () can contribute to LVK detections (Kinugawa et al. 2014). Assuming a metallicity range for LVK we obtain a broad range of BH mass: . This range is consistent with the LVK BH mass estimates in BH-BH mergers. The recent discovery of a BH in the binary Gaia BH3 (Gaia Collaboration et al. 2024) strongly supports the scenario according to which metal-poor massive stars are progenitors of the high-mass BHs observed by LVK (see, e.g. Olejak et al. 2020).
In conclusion, the same model of stellar evolution produces correct mass ranges for the three discussed HMXB BHs (narrow) and LVK BHs (wide).
@@ 3 BH spins: stellar evolution
Population synthesis models are also used to predict the BH spin distribution. In most cases, they produce low-spinning BHs () from non-interacting stars and low- to rapidly-spinning BHs () for Wolf-Rayet (WR) stars that were subject to tidal spin-up in close binaries (Belczynski et al. 2020b; Bavera et al. 2020). Such models can explain the low and high effective spins of LVK BH-BH mergers (Olejak & Belczynski 2021).
@@
The low BH spins obtained by population synthesis models are the consequence of the assumed efficient angular momentum transport from stellar cores to envelopes, an effect that has been now well confirmed by observations (Aerts et al. 2019; Langer 2012) consistent with the action of the magnetic Tayler (1973) instability (Spruit 2002a; Fuller et al. 2019; Eggenberger et al. 2022; Petitdemange et al. 2024). The heuristic Spruit (2002a) version of the mechanism used in population synthesis codes has been recently confirmed by numerical calculations of Petitdemange et al. (2024).
@@
Therefore unless their progenitors in HMXBs belong to a special category of stars with uncoupled cores and envelopes, BHs in these systems should be formed with low spins (). The known HMXB orbital periods are too long for tidal interactions to spin up the stars before they form BHs (the size of the present BH’s companion would not fit into a days orbit of the tidal locking regime; see e.g. Belczynski et al. 2020a) so if their spins are really large they must have have gained angular momentum by accretion. To spin up a BH from to one has roughly to double its mass (Bardeen 1970). For Cyg X-1 this would require accreting in less than Myr (Miller-Jones et al. 2021), i.e an accretion rate , or higher by factor of 100 if the accretion lifetime of Cyg X-1 is only a few tens of thousands of years Miller-Jones et al. 2021). This gives an accretion rate of where is the Eddington accretion rate with an assumed efficiency of . Contrary to the still expressed opinion (see e.g. Miller-Jones et al. 2021), the Eddington rate is not a maximum theoretical limit (Begelman 1979) and black holes can accrete from discs at rates up to at least a few tens of (Kitaki et al. 2021; Yoshioka et al. 2022, but see Hu et al. 2022), so that, an accretion rate in Cyg X-1 is not, from this point of view, extravagant. The problem is to find an evolutionary path that would take a high-mass binary through such a super-Eddington phase and produce the observed Cyg X-1, whose O supergiant donor star has yet to fill its Roche lobe. This is unlikely to happen for accretion from wind, even when it is L1-focused (Wiktorowicz et al. 2021). Therefore, according to the best stellar physics available, BH spins in the known HMXBs should be low, with – consistent with LVK BH spins (Fishbach & Kalogera 2022) and estimates of BH spin in HMXBs by Belczynski et al. (2021); Zdziarski et al. (2024a, b), but clearly incompatible with the generally accepted high spin values in these systems (Miller & Miller 2015). These high spin values have been obtained by assuming the presence of a standard disc in the system (see Sect. 7 for the standard-disc definition), which can be justified only if the system is in a soft spectral state. In the following we will argue that since Cyg X-1 has never been observed in such a state, the high black hole spin from disc continuum fitting is not a firm requirement.
@@ 4 BH spins from disc spectral fitting: HMXBs
Standard accretion disc models balance gravitational heating from viscous torques with blackbody cooling, resulting in an optically thick, geometrically thin structure with effective temperature at each radius, . This gives a total spectrum, which is a sum of all radii of blackbody spectra of different temperatures. The emissivity peaks around , so the peak temperature is K, i.e. is a factor higher for the same disc luminosity at high spin than at low spin.
In all X-ray binaries, one observes cycles during which the brightness and spectral hardness form a characteristic “tortoise-head” pattern. In general, the spectrum contains both thermal disc emission, together with a power law tail extending to much higher energies, clearly showing that there is an additional non-disc component to the accretion flow. Most transients show a phase in this cycle where the spectrum is sufficiently soft to be termed disc dominated, where the assumption of accretion disc assumption is likely to be reliable. The problem is that not all X-ray sources complete the full cycle. As we will show below, this is the case for Cyg X-1.
Indeed, there is clear evidence that in Cyg X-1, the accretion flow has not formed a standard black body disc. The blue points in Figure 2 show the hardness-intensity diagram of the LMXB GX339-4. This source picks out a well-defined track, transitioning from low/hard (lower right) to disc-dominated (upper left, hardness ratio below ), with fast transitions (horizontal) between the two states. The red points compare data from Cyg X-1 (both data taken from Belloni 2010). The mass and distance of GX339-4 are not well known, so we have scaled the relative count rate (brightness) between the two systems to match the lower transition luminosity. We can see how GX339-4 transitions from low/hard state (hardness ratio ; the ratio is defined as the ratio of counts in the energy bands and keV) to high/soft state (hardness ratio ). In contrast, Cyg X-1 never reaches as far to the left as seen in the GX339-4 disc-dominated states.
Hence, Cyg X-1 always has a strong tail of emission to higher energies (e.g. Gierliński et al. 1999; Gou et al. 2009; Walton et al. 2016), making it stuck in the more complex intermediate states where there is no reason to believe that the accretion flow forms a standard accretion disc. Instead, the soft component in intermediate state spectra is often better fit by optically thick, warm (few keV) Comptonization (see, for e.g. transition spectra from the LMXB-BH MAXI J1820+070 in Fig. 2 of Kawamura et al. 2023).
@@
Instead, the strong soft component need not be emitted by a simple blackbody disc, where the energy is dissipated mainly on the midplane but can have heating further up into the photosphere. This forms a warm layer on the top of the disc, making it appear hotter and giving higher spin when fit with standard disc models. Such a warm layer on the disc is sometimes required in the black hole binaries (e.g. in very high/intermediate type spectra: Kubota et al. 2001; Remillard & McClintock 2006) and is generic in Active Galactic Nuclei (Porquet et al. 2019; Gierliński & Done 2004b; Petrucci et al. 2018). On the other hand it is well known that the simplistic assumption of dissipation concentrated at the disc midplane (usually in the form of dissipation proportional to pressure) is in contradiction with both observations (see e.g., Hubeny & Long 2021) and numerical simulations (Blaes 2014).
Thus, there is no physical reason to assume that a standard accretion disc emits the observed X-ray spectrum of Cyg X-1. In fact, there are good reasons to think that such an assumption is inappropriate for this system. There is no “canonical” model describing the inner disc structure in X-ray binaries – no model correctly describes the “tortoise-head” cycle. Hence, one is allowed to test various options that would reproduce the observed spectrum and check if the deduced spin value is model-dependent, and if it is, which model corresponds best to the value obtained from stellar evolution. A similar approach was adopted by Zdziarski et al. (2024a, b).

### v4

# Common origin for black holes in both high mass X-ray binaries and gravitational-wave sources

@@
Common origin for black holes in both high mass X-ray binaries and gravitational-wave sources
Black-hole (BH) high-mass X-ray binary (HMXB) systems are likely to be the progenitors of BH-BH mergers detected in gravitational waves by LIGO/Virgo/KAGRA (LVK). Yet merging BHs reach higher masses () than BHs in HMXBs () and typically exhibit lower spins ( with a larger values tail) than what is often claimed for BHs in HMXBs (). This could suggest that these two classes of systems belong to different populations, but here we show that this may not necessarily be the case. The difference in masses is easily explained as the known HMXB-BHs are in galaxies with relatively high metallicity, so their progenitor stars are subject to strong mass loss from winds, leading to relatively low-mass BH at core collapse. Conversely, LVK is also able to detect BHs from low-metallicity galaxies that are known to naturally produce more massive stellar-origin BHs. However, the difference in spin is more difficult to explain. Models with efficient angular momentum transport in stellar interiors produce slowly spinning progenitors for both LVK and HMXB BHs. Known HMXBs have orbital periods that are too long for efficient tidal spin-up and are also unlikely to have undergone significant accretion spin-up. Instead, we show that the derived value of the BH spin depends strongly on how the HMXB accretion disc emission is modelled. We argue that since Cyg X-1 is never observed to be in a soft spectral state, the appropriate spectral models must take into account the Comptonisation of the disc photosphere. We show that such models are consistent with low spin values, namely: . This was recently confirmed by other teams for both Cyg X-1 and LMC X-1 and here we show this is also the case for M33 X-7. We conclude that all known HMXB BHs can exhibit a low spin, in accordance with the results of stellar evolution models. Hence, the observations presented in this work are consistent with the LVK BHs and HMXB BHs belonging to the same population.
@@ 1 Introduction
LIGO/Virgo/KAGRA (LVK) interferometers have detected gravitational waves from double black hole (BH-BH) mergers (The LIGO Scientific Collaboration et al. 2023). These merging black holes (BHs) have masses in the range , with many primary (the more massive) BHs having masses of and (Callister & Farr 2024). The most massive LVK event (GW190521) showed two merging BHs with estimated mass of and . The majority of these mergers have low positive effective spin parameters, with the distribution peaking at
@@
where denotes BH masses, are the BH spin magnitudes (: BH angular momentum, with as the speed of light and as the gravitational constant), while refers to the angles between the system’s angular momentum and BH spins. Out of these mergers, of them show high positive effective spins (mean), while none show a high negative spin . However, there are three events with moderate negative spin estimates: . Low values of the effective spin parameter may result from small individual BH spin magnitudes (Belczynski et al. 2020b). This is supported by LVK data, which shows that BH individual spin magnitudes in BH-BH mergers peak at ; however, the distribution has a long tail that extends to large values (The LIGO Scientific Collaboration et al. 2023; see Callister & Farr 2024 for slightly different results).
@@
Electromagnetic observations have revealed a population of BHs hosted in binary star systems111e.g. see https://universeathome.pl/universe/blackholes.php and references therein. Out of several known binary configurations hosting BHs, only high-mass X-ray binaries (HMXBs) can potentially lead to BH-BH mergers, as BH companion stars are massive enough to form a BH. There are only three HMXBs known to have such massive () companions and they have moderate-mass BHs: LMC X-1 with a star and a BH (Orosz et al. 2009); Cyg X-1 with a star and a BH (Miller-Jones et al. 2021); and M33 X-7 with a star and a BH (Ramachandran et al. 2022; revised down from the previous estimates of a donor and BH in Orosz et al. 2007). We note that the binary LMC X-3 is not a HMBXB since its stellar companion is less massive () than the BH () Orosz et al. (2014) and will not form a BH anyway. The spins of BHs in HMXBs were obtained with two methods, namely: disc continuum fitting, and reflection spectroscopy. Reflection spectroscopy is complex for soft state spectra (e.g. Tomsick et al. 2018), so here we focus on spin values from a disc continuum fitting. These are typically very high222The spins attributed through various methods to LMXB BH vary from very low to very high values, sometimes for the same system (e.g. Reynolds 2021). However, on the one hand these binary systems are not BHBH binary progenitors, on the other, their evolutionary history is still a puzzle (Wiktorowicz et al. 2014).:
@@
Since masses and spins of LVK BHs and HMXB BHs appear to be different, it might seem that these two groups of systems belong to populations with different formation scenarios. Here, we show that all these BHs may instead form a single population. The difference in BH mass arises naturally from the diverse formation environments: all known HMXB-BH systems reside in Local Group galaxies with relatively high metallicity, so they have strong stellar winds before the core collapse. This results in lower black hole masses compared to those forming in low metallicity galaxies, which are within the sensitivity range of LVK. We demonstrate the difference in expected black hole masses explicitly using a population synthesis model. This model also reproduces the spin values in the LVK BHs, for models where high spins are produced only by tidal spin-ups in extremely close binaries. The known HMXB-BH systems are in binary systems that are not close enough for tidal spin-up. Hence, the population synthesis models predict they should not have high spins, which conflicts with the reported high values. However, the disc continuum method relies on the emission thermalising. If, instead, the disc is covered by a warm, optically thick skin, whereby the emission is Comptonised, the spin values are all consistent with (Belczynski et al. 2021; Zdziarski et al. 2024a, b; this paper). With this interpretation, the LVK BH-BH mergers and the HMXBs can be considered as members of the same binary star population.
@@ 2 BH masses
To demonstrate the unicity of the BH population, we use the population synthesis code StarTrack (Belczynski et al. 2002, 2008) and employ the delayed core-collapse supernova (SN) engine for neutron star/BH mass calculation (Fryer et al. 2012) that does not produce a lower mass gap between NSs and BHs. We used a model in which the pair-instability (upper) mass gap is calculated with non-standard fusion reaction rates and heavy mixing for massive stars; this allows for BHs to form up to (Farmer et al. 2020; Costa et al. 2021). This type of model had already been developed by Belczynski (2020) in the context of the detected at the low-metallicity GW190521 LVK merger, at distance of Gpc (Abbott et al. 2020).
In Fig. 1, we show the dependence of the helium core mass on metallicity for single stars obtained with the stellar wind treatment presented in Belczynski et al. (2010a). The helium core mass is a good approximation of the BH mass since binary interactions (Roche Lobe overflow or common envelope evolution) tend to remove the H-rich envelope from stars in close binaries (progenitors of BH-BH mergers if these form through isolated binary evolution). The lower the metallicity, the lower the opacity in stellar atmospheres, and the lower the wind mass loss, which produces progressively higher mass BHs with decreasing metallicity.
Cyg X-1 is located in the Milky Way disc, which hosts stars that have metallicity values that are close to solar (, Asplund et al. 2009). For such values, we obtain a maximum BH mass of , which is consistent with the highest mass stellar-origin BH known in the Milky Way for the same Cyg X-1 (, Miller-Jones et al. 2021). LMC X-1 is located in the Large Magellanic Cloud () and M33 X-7 in M33 galaxy (). In this metallicity range, BHs are not expected to be more massive than .
LVK BH-BH mergers have already been detected (O3) to significant redshifts of . The detected BH-BH mergers may have formed even at much larger redshifts due to non-zero delay times between star formation and BH-BH mergers (Dominik et al. 2012; Fishbach & Kalogera 2021). Therefore, galaxies and stars that can produce BH-BH mergers cover a wide range of metallicities. It is even claimed that Population III (metal-free) stars () may contribute to LVK detections (Kinugawa et al. 2014). Assuming a metallicity range for LVK we obtain a broad range of BH mass: . This range is consistent with the LVK BH mass estimates in BH-BH mergers. The recent discovery of a BH in the binary Gaia BH3 (Gaia Collaboration et al. 2024) strongly supports a scenario in which metal-poor massive stars are progenitors of the high-mass BHs observed by LVK (see e.g. Olejak et al. 2020). In conclusion, the same model of stellar evolution produces correct mass ranges for the three discussed HMXB BHs (narrow) and LVK BHs (wide).
@@ 3 BH spins: Stellar evolution
Population synthesis models are also used to predict the BH spin distribution. In most cases, they produce low-spinning BHs () from non-interacting stars and slow-to-rapidly-spinning BHs () for Wolf-Rayet (WR) stars that were subject to tidal spin-up in close binaries (Belczynski et al. 2020b; Bavera et al. 2020). Such models can explain the low and high effective spins of LVK BH-BH mergers (Olejak & Belczynski 2021).
@@
The low BH spins obtained by population synthesis models are the consequence of the assumed efficient angular momentum transport from stellar cores to envelopes, an effect that has been now aptly confirmed by observations (Aerts et al. 2019; Langer 2012) consistent with the action of the magnetic Tayler (1973) instability (Spruit 2002a; Fuller et al. 2019; Eggenberger et al. 2022; Petitdemange et al. 2024). The heuristic Spruit (2002a) version of the mechanism used in population synthesis codes was recently confirmed by numerical calculations of Petitdemange et al. (2024).
@@
Therefore, unless their progenitors in HMXBs belong to a special category of stars with uncoupled cores and envelopes, the BHs in these systems would have been formed with low spins (). The known HMXB orbital periods are too long for tidal interactions to spin up the stars before they form BHs (the size of the present BH companion would not fit into a days orbit of the tidal locking regime; see e.g. Belczynski et al. 2020a); thus, if their spins are really large, they must have have gained angular momentum by accretion. To spin-up a BH from to it is necessary to roughly to double its mass (Bardeen 1970). For Cyg X-1, this would require accreting in less than Myr (Miller-Jones et al. 2021), namely, with an accretion rate of ; this rate would be higher by factor of 100 if the accretion lifetime of Cyg X-1 was only a few tens of thousands of years (Miller-Jones et al. 2021). This gives an accretion rate of , where is the Eddington accretion rate with an assumed efficiency of . Contrary to the still expressed opinion (see e.g. Miller-Jones et al. 2021), the Eddington rate is not a maximum theoretical limit (Begelman 1979) and black holes can accrete from discs at rates of up to (at least) a few tens of (Kitaki et al. 2021; Yoshioka et al. 2022, but see Hu et al. 2022); thus, an accretion rate in Cyg X-1 would not be, from this point of view, extravagant. The problem is to find an evolutionary path that would take a high-mass binary through such a super-Eddington phase and produce the observed Cyg X-1, whose O supergiant donor star has yet to fill its Roche lobe. This is unlikely to happen with respect to accretion from wind, even when it is L1-focused (Wiktorowicz et al. 2021). Therefore, according to the best stellar physics available BH spins in the known HMXBs should be low, with – consistent with LVK BH spins (Fishbach & Kalogera 2022) and estimates of BH spin in HMXBs by Belczynski et al. (2021); Zdziarski et al. (2024a, b); however, this is clearly incompatible with the generally accepted high spin values in these systems (Miller & Miller 2015). These high spin values have been obtained by assuming the presence of a standard disc in the system (see Sect. 7 for the standard-disc definition), which can be justified only if the system is in a soft spectral state. In the following, we argue that since Cyg X-1 has never been observed in such a state, the high black hole spin from disc continuum fitting is not a firm requirement.
@@ 4 BH spins from disc spectral fitting: HMXBs
Standard accretion disc models balance gravitational heating from viscous torques with blackbody cooling, resulting in an optically thick, geometrically thin structure with effective temperature at each radius, . This gives a total spectrum, which is a sum of all radii of blackbody spectra of different temperatures. The emissivity peaks around , so the peak temperature is K, namely, it is a factor higher for the same disc luminosity at high spin than at low spin.
In all X-ray binaries, we can observe cycles during which the brightness and spectral hardness form a characteristic ‘tortoise-head’ pattern. In general, the spectrum contains both thermal disc emission, together with a power law tail extending to much higher energies, clearly showing that there is an additional non-disc component to the accretion flow. Most transients show a phase in this cycle where the spectrum is sufficiently soft to be termed ‘disc-dominated’, where the accretion disc assumption is likely to be reliable. The problem is that not all X-ray sources complete the full cycle. As we show below, this is the case for Cyg X-1.
Indeed, there is clear evidence that in Cyg X-1, the accretion flow has not formed a standard blackbody disc. The blue points in Fig. 2 show the hardness-intensity diagram of the LMXB GX339-4. This source picks out a well-defined track, transitioning from low and hard (bottom-right) to disc-dominated (upper left, hardness ratio below ), with fast transitions (horizontal) between the two states. The red points compare data from Cyg X-1 (both data taken from Belloni 2010). The mass and distance of GX339-4 are not well known, so we have scaled the relative count rate (brightness) between the two systems to match the lower transition luminosity. We can see how GX339-4 transitions from low/hard state (hardness ratio ; the ratio is defined as the ratio of counts in the energy bands and keV) to high/soft state (hardness ratio of ). In contrast, Cyg X-1 never reaches as far to the left as seen in the GX339-4 disc-dominated states.
Hence, Cyg X-1 always has a strong tail of emission to higher energies (e.g. Gierliński et al. 1999; Gou et al. 2009; Walton et al. 2016), keeping it stuck in the more complex intermediate states where there is no reason to believe that the accretion flow forms a standard accretion disc. Instead, the soft component in intermediate state spectra is often better fit by optically thick, warm (few keV) Comptonisation (see e.g. transition spectra from LMXB-BH MAXI J1820+070 in Fig. 2 of Kawamura et al. 2023).
@@
Instead, the strong soft component need not be emitted by a simple blackbody disc, where the energy is dissipated mainly on the midplane but can have heating further up into the photosphere. This forms a warm layer on the top of the disc, making it appear hotter and giving a higher spin when fit with standard disc models. Such a warm layer on the disc is sometimes required in the black hole binaries (e.g. in very high and intermediate type spectra: Kubota et al. 2001; Remillard & McClintock 2006) and is generic in active galactic nuclei (Porquet et al. 2019; Gierliński & Done 2004b; Petrucci et al. 2018). On the other hand, it is well known that the simplistic assumption of dissipation that is concentrated at the disc midplane (usually in the form of dissipation proportional to pressure) is in contradiction with both observations (see e.g. Hubeny & Long 2021) and numerical simulations (Blaes 2014).
Thus, there is no physical reason to assume that a standard accretion disc emits the observed X-ray spectrum of Cyg X-1. In fact, there are good reasons to think that such an assumption is inappropriate for this system. There is no ‘canonical’ model describing the inner disc structure in X-ray binaries – no model correctly describes the ‘tortoise-head’ cycle. Hence, we can test various options that would reproduce the observed spectrum and check whether the deduced spin value is model-dependent; if it is, we consider which model corresponds best to the value obtained from stellar evolution. A similar approach was adopted by Zdziarski et al. (2024a, b).